OpenStudy (anonymous):

At East Zone University (EZU) there are 655 students taking College Algebra or Calculus. 293 are taking College Algebra, 384 are taking Calculus, and 22 are taking both College Algebra and Calculus. How many are taking Calculus but not Algebra?

OpenStudy (blockcolder):

Let A=the set of students taking Algebra and C=the set of students taking calculus. We're looking for |C-A|. \[|C-A|=|C|-|C\cap A|\] \[=384-22=362.\]

OpenStudy (anonymous):